-[5r-(8r+6)]=6(r-5)-12
Answer by stanbon(75887)

-[5r-(8r+6)] = 6(r-5)-12
-5r + (8r+6) = 6r - 30 -12
3r + 6 = 6r - 42
3r = 48
r = 16
--------------------------
Cheers,
Stan H.